OpenStudy (anonymous):

aRb (a less or equal to (-b)) and it consist of all integers. is it symmetric,reflexive, and transitive. oh man I hate relations And I bet my answers are all wrong ,but I just wanted to check :(

OpenStudy (helder_edwin):

so u have \[\large (\forall a,b\in\mathbb{Z})(aRb\Leftrightarrow a\leq-b) \] right?

OpenStudy (anonymous):

yup

OpenStudy (helder_edwin):

do u know the definitions?

OpenStudy (anonymous):

what definitions ? I just want the answers , since I just had a mid term exam and this one was the last question.

OpenStudy (helder_edwin):

ok

OpenStudy (helder_edwin):

do u think that \[\large a\leq-a \] for every integer a?

OpenStudy (anonymous):

it's not reflexive for what I know. anyways I said that they're all true since he didn't give me the quantifiers for the values of x and y , it just said that they're integers , after the exam I thought of it and I think we should work on every integer not just give a counter example of some integer

OpenStudy (helder_edwin):

well, in this the quantifiers were implicit (they usually are)

OpenStudy (anonymous):

Our prof only gives one example of each topic and I remember him using a counter example to solve most of the examples. anyways I think that they're all false since we're talking about all integers.

OpenStudy (helder_edwin):

counterexamples are used to show something is not true when the quantifier is \(\forall\)

OpenStudy (anonymous):

:( anyways are they all false or true ?

OpenStudy (helder_edwin):

in this case the relation IS symmetric

OpenStudy (helder_edwin):

if aRb then \[\large a\leq -b \] multiplying by (-1) we get \[\large b\leq -a \] which mean bRa. So R is symmetric

OpenStudy (anonymous):

Even though I said it's symmetric , I think I won't get any marks since he wants a full explanations or he'll simply consider it false and give a 0 for it.

OpenStudy (anonymous):

so it's only symmetric ?

OpenStudy (helder_edwin):

it is not transitive because \[\large 1R(-2)\quad\text{because}\quad1\leq2 \] and \[\large (-2)R0\quad\text{because}\quad-2\leq0 \] but 1R0 is false because \[\large 1\not\leq0 \]

OpenStudy (anonymous):

Thank you :)

OpenStudy (helder_edwin):

u r welcome